The length of a rectangle is twice the width. The area is 18. Find the dimensions of the rectangle.
Answer by checkley77(12844)

L=2W
LW=18
2W*W=18
2W^2=18
W^2=18/2
W^2=9
W=SQRT9
W=3 ANS. FOR THE WIDTH.
L=2*3
L=6 ANS. FOR THE LRNGTH.
PROOF:
6*3=18
18=18